OpenStudy (anonymous):

What is the equation of the line that passes through the points (2, –1) and (6, 1)?

OpenStudy (itiaax):

First, find the graient of the line..do you know how to find it?

OpenStudy (itiaax):

gradient*

OpenStudy (anonymous):

no

OpenStudy (itiaax):

The gradient is found by using this formula: y2-y1/x2-x1 where x and y are the coordinates. So, the gradient will be: (1--1)/(6-2) and we end up with 2/4 which can be simplified to 1/2

OpenStudy (itiaax):

Then, inorder to find the equation of the line, we use the formula y-mx+c, where y is y-coordinate, m is the gradient/slope of the line, x is the x-coordinate and c is the y-intercept (the place where the line cuts the y-axis). So, let's say we use the coordinates (6,1), substituting, we get: 1=1/2(6)+c. We have to solve for c, so we get 1=3+c. c=1-3 = -2. Thus, the equation of the line is y=1/2x-3

OpenStudy (anonymous):

so its y=-2x-3 or y=-1/2x-1

OpenStudy (anonymous):

if this is a straight line, use y=mx+b. first find the slope: m=y2-y1/x2-x1 =1+1/6-2 =2/4 slope=1/2 now plug in any point with the slope to find b y=mx+b (1)=(1/2)(6)+b 1=6/2+b 1=3+b -2=b so equation is: (y=mx+b) y=1/2 x -2